For a gas compressor compresses gas (Cv= 718 J/KgK and Cp= 1005 J/KgK) adiabatically from 1 bar and 15 ºC to 10 bar with an isentropic efficiency of 0.89. The gas flow rate is 5 kg/s. Calculate the temperature after compression and the power input
Answer #1

answer) for temperature after compression we use

T2=T1*(P2/P1)1-1/

=cp/cv=1005/718=1.4

T1=273+15=288 K

T2=288K*(10/1)1-1/1.4

T2=556 K

so temperature=556 K ( answer)

power input=actual power/efficiency

actual power=m*cp*dt=5*1005*(556-288)=1.35*106 W

power input=1.35*106 W/0.89=1.51*106 W

so the answer is 1.51 MW or 1.51*106 W.